OpenStudy (anonymous):

HELP PLEASE! :) Find the surface area of an orange with a 2'' radius.

OpenStudy (amistre64):

not possible with the given information :)

OpenStudy (anonymous):

We can assume, i hope, that the orange is a sphere ?

OpenStudy (anonymous):

one second i forgot somehting

OpenStudy (amistre64):

there needs to be alot of assumption in order to make this work right ;)

OpenStudy (anonymous):

Use pie == 22/7

OpenStudy (anonymous):

\[\A = 4\pi r^2\]

OpenStudy (anonymous):

\[4*(22/7)*2^2\]

OpenStudy (anonymous):

A spherical storage tank has a diameter of 14 ft. How many cubic feet of water will it hold? (Use pie = 22/7

OpenStudy (anonymous):

V=(4/3)*(pi)*(r)^3

OpenStudy (anonymous):

plug in 22/7 for pi and 7 for the radius, half of the diameter

OpenStudy (anonymous):

V=(4/3)*(22/7)*(7)^3

OpenStudy (anonymous):

multiply it?

OpenStudy (anonymous):

\[(4/3)*(22/7)*343\]

OpenStudy (anonymous):

yeah multiply

OpenStudy (anonymous):

um im not typing thats weird

OpenStudy (anonymous):

\[(4/3)*22*49\]

OpenStudy (anonymous):

?

OpenStudy (anonymous):

12936?

OpenStudy (anonymous):

nope

OpenStudy (anonymous):

did i multipy wrong

OpenStudy (anonymous):

you cant multiply?lol

OpenStudy (anonymous):

\[4312/3\]

OpenStudy (anonymous):

1437.3333333 etc

OpenStudy (anonymous):

yea :D

OpenStudy (anonymous):

i think i have to round it

OpenStudy (anonymous):

thank you :D

OpenStudy (anonymous):

1437.33